OpenStudy (anonymous):

Solve the inequality. Write your answer using interval notation. Please give me a second to write up the inequality

OpenStudy (anonymous):

\[\frac{ 2x-3 }{ x+1 }\le1\]

OpenStudy (anonymous):

my first thought is to make x the subject of the formula but ik that in incorrect and ill end up with only one answer. I cannt remember the method to get two answers so id be nice if you can explain as your solve the question.

jimthompson5910 (jim_thompson5910):

what do you get when you solve the equation below for x (2x-3)/(x+1) = 1

OpenStudy (anonymous):

i would do it like this 2x-3 = x+1 and make x the subject of the formula

OpenStudy (anonymous):

and get x = 4

OpenStudy (anonymous):

first solve for x

jimthompson5910 (jim_thompson5910):

good, now what is the vertical asymptote of (2x-3)/(x+1)

OpenStudy (anonymous):

vertical asymptote?

OpenStudy (anonymous):

 plot 4 on this type of graph